Find the position and velocity of a particle at t = 1.98 s if the particle is initially moving east at a speed of 19.8 m/s and experiences an acceleration of magnitude 4.40 m/s2, directed west.
Answer #1

u=19.8

a=-4.4

v=u+at=19.8-4.4*1.98=11.088 m/s

s=ut+0.5at^2=19.8 *1.98-0.5*4.4*1.98^2=30.57
Answer #2

v = u + at = 19.8 - 4.40 x 1.98 = 11.09 m/s

s = ut + at^2 /2 = 19.8 x 1.98 + (-4.40|) x 1.98^2 /2 = 30.59 m